Ship reports banner with links Home IODP-USIO daily/weekly reports

< Previous expedition | January >

IODP Expedition 345

Hess Deep Plutonic Crust

Daily Science Report for 31 December 2012

LOCATION: Site U1415 (HD-01B), Hess Deep

SCIENCE UPDATE: Today, we assembled a 14¾ inch tricone bit, lowered it to the seafloor, and lowered the camera system to search for Hole U1415I. We easily found the mound of cuttings around Hole U1415I. Although we could see where the hole had been, we could not identify that the hole was open. It appeared that rubble had collapsed into the top of the hole. We decided to abandon Hole U1415I. We offset the ship 10 m to the north, checked the seafloor depth with the bit, retrieved the camera system, and initiated Hole U1415J at 1855 h. Our plan is to use this 14.75 inch hole to isolate the uppermost unstable formation with casing.


Daily Science Report for 30 December 2012

LOCATION: Site U1415 (HD-01B), Hess Deep

SCIENCE UPDATE: Coring continued today and Core U1415I-4R arrived on deck just before 0800 h. It was cut from 27.0 to 35.2 m and recovered 1.6 m of very nice gabbroic rocks (20% recovery). We dropped the next core barrel but encountered hole problems attempting to get back down to the bottom of the hole. After spending a couple hours washing and reaming between 7.7 and 11.0 mbsf, we decided that we would trip out of the hole to change bits. In preparation to deploy a free-fall funnel (FFF), we deployed the camera system to observe the seafloor to determine if dropping a FFF would be feasible. While doing this, the hole conditions continued to deteriorate and we were unable to penetrate past 5 mbsf. Based on this we decided not to deploy a FFF and we pulled out of the hole. We retrieved the core barrel (Core U1415I–5G) that contained 21 cm of gabbroic rocks from the upper portion of the hole. The latter part of the day was spent recovering the drill string. We plan to assemble a 14¾ inch tricone bit and either open up the existing hole or start a new one. This may allow us to isolate the uppermost unstable formation with casing.


Daily Science Report for 29 December 2012

LOCATION: Site U1415 (HD-01B), Hess Deep

SCIENCE UPDATE: Cores U1415HI-1R to 3R extend from the seafloor to 27 mbsf. The cores recovered ~0.92 m of gabbroic rocks. We also recovered ~4.6 m of coarse sand in the top of Core 3R, but this may be drill cuttings that we are having trouble clearing from the hole. The drilling continues to advance, but hole instability (rubble?) is causing periodic problems. In addition, the hole periodically packs off (indicated by increased pump pressures) the suggests to the drillers that we are not able to fully circulate the cuttings out of the hole.


Daily Science Report for 28 December 2012

LOCATION: Site U1415 (HD-01B), Hess Deep

SCIENCE UPDATE: Today started with RCB coring in Hole U1415H. Core U1415H-1R penetrated to 12.9 mbsf and recovered 0.44 m (4%) of gabbroic rocks. After raising the bit off the bottom to retrieve the core, the hole conditions were similar to previous holes where we had significant trouble and damaged bottom-hole assemblies. So we terminated coring in Hole U1415H and pulled the bit out of the hole. We moved the ship ~50 m to the northwest, inspected the seafloor with the camera system, and verified the seafloor depth with the bit. After retrieving the camera system, we started coring in Hole U1415I. By the end of the day, RCB coring had reached 6.5 mbsf; drilling conditions were consistent, reasonably good, but with slow penetration. Scientists and technical staff attended a presentation on the IODP Superfast Expedition Hole 1256D.


Daily Science Report for 27 December 2012

LOCATION: Site U1415 (HD-01B), Hess Deep

SCIENCE UPDATE: Today started with RCB coring in Hole U1415G. Core U1415G-1R penetrated to 11.4 mbsf and recovered 0.29 m (3%) of gabbroic rocks. After raising the bit off the bottom to retrieve the core, the hole conditions were similar to previous holes where we had significant trouble and damaged bottom-hole assemblies. So we terminated coring in Hole U1415G and pulled the bit out of the hole. We moved the ship (20 m west; 15 m south), verified the seafloor depth with the bit, and then started coring in Hole U1415H. By the end of the day, we had cored to 8.2 mbsf and drilling conditions were consistent and reasonably good. Scientists and technical staff attended a presentation the formation of sheeted dike complexes at fast spreading centers through variable magma flow directions.


Daily Science Report for 26 December 2012

LOCATION: Site U1415 (HD-01B), Hess Deep

SCIENCE UPDATE: We continued retrieving the drill string from Hole U1415E. The bottom-hole assembly (BHA) had failed at the bottom of the fourth drill collar and two drill collars above this were bent. We assembled a new RCB BHA, lowered it to the seafloor, and conducted an ~4 hr survey with the camera system and 3.5 kHz sub-bottom pinger. At the end of the day, we conducted a jet-test that reached 1.5 mbsf (Hole U1415F) and then prepared to start RCB coring in Hole U1415G. Scientists and technical staff attended a presentation on creating the lithosphere at Gakkel Ridge in the Arctic Ocean.


Daily Science Report for 25 December 2012

LOCATION: Site U1415 (HD-01B), Hess Deep

SCIENCE UPDATE: Today started with a brief jet-in test that indicated a little over 4 m of sediment before the bit encountered hard formation (Hole U1415D). We pulled out of the seafloor, retrieved the camera system, and then started RCB-coring in Hole U1415E without offsetting the ship. This time a hard formation was encountered at ~2 mbsf. Cores U1415E-1R and -2R were extended from 0 to 15.3 mbsf and recovered 0.84 m of gabbroic rocks (5%). Drilling conditions were poor. We observed erratic drill string torque and when we had to pick up off bottom to make a connection the hole collapsed and we had to fight our way back down to the bottom of the hole. After recovering Core 2R, we observed a drop in pump pressure, a reduction in torque, and a drop in drill string weight. We inferred that the bottom-hole assembly had failed and spent the rest of the day retrieving the drill string.


Daily Science Report for 24 December 2012

LOCATION: Site U1415 (HD-01B), Hess Deep

SCIENCE UPDATE: We finished retrieving the drill string and had to carefully check all parts of the bottom-hole assembly (BHA). In the end, much of the top parts of the BHA were damaged (bent) and had to be retired. This included five drill collars, the tapered drill collar and one joint of 5.5 inch drill pipe. After assembling a new RCB BHA, we lowered it to the seafloor. Once the camera system with the 3.5 kHz pinger reached the seafloor, we started another near-bottom 3.5 kHz survey to assess sediment thickness. Unfortunately, the pinger stopped transmitting shortly after starting and the sub-bottom survey was terminated. We kept the camera system down and moved a short distance (~100 m) to the next location. At the end of the day, we had conducted a short camera survey around the location and initiated a jet-in test (Hole U1415D).


Daily Science Report for 23 December 2012

LOCATION: Site U1415 (HD-01B), Hess Deep

SCIENCE UPDATE: Today started with retrieving the camera system to turn the 3.5 kHz pinger on. While the camera system was on board, we verified the seafloor depth at proposed site HD-01B with the bit (4848.5 mbrf). We then conducted a jet-in test that reached a depth of 10.3 mbsf (Hole U1415A). After the jet-in test, we lowered the camera system with the attached 3.5 kHz pinger and conducted a near-bottom 3.5 kHz survey to assess sediment thickness. We conducted two more jet-in tests at a single location later in the day. The first reached 11.7 m (Hole U1415B), but we needed to adjust our pipe space out before continuing deeper, so we raised the bit above the seafloor. The second apparently reach 76.5 mbsf (Hole U1415C). During this test, we retrieved the camera system. Once it was on board, we attempted to rotate the drill string, but could not until we raised the bit above the seafloor. We inferred that the bottom-hole assembly may have been bent, so the rest of the day was spent recovering the drill string. Scientists attended a presentation on the modeling the diffusion of magnesium in plagioclase to determine cooling rates of the lower oceanic crust.


Daily Science Report for 22 December 2012

LOCATION: Hess Deep proposed site HD-01B

SCIENCE UPDATE: The second half of the science party (metamorphic, structure, and geochemistry teams) attended the Curator​s introduction to core flow and sampling. After arriving on site just before midnight last night, we started assembling the RCB bottom-hole assembly and lowering it to the seafloor, which is at ~4855 m below sea level. After the bit was near bottom along with the camera system, we conducted a short survey to verify the seafloor topography where our sites are located, deployed a seafloor positioning beacon from the camera frame ~100 m to the west of proposed site HD-01B, and then prepared for a jet-in test at site HD-01B.


Daily Science Report for 21 December 2012

LOCATION: In Transit to Hess Deep, arrived site HD-01B at 2350 h

SCIENCE UPDATE: The lab teams continued preparing their methods, instruments, and workflow, facilitated by using ODP Leg 147 Hess Deep cores, samples, and data. The laboratory team submitted their first draft of their methods for review. Scientists and technical staff attended a presentation on deep magmatic processes beneath fast-spreading ridges. A Chief Scientist then provided an update on the plans for the pre-drilling survey of sediment thickness as well as initial drilling/​coring operations. Half of the science party (igneous, physical properties, and paleomagnetism teams) attended the Curator’s introduction to core flow and sampling. Just before midnight (2350 h) we arrived at Hess Deep proposed site HD-01B, started lowering the thrusters and hydrophones, and began to position the ship over the drill site.


Daily Science Report for 20 December 2012

LOCATION: In Transit to Hess Deep

SCIENCE UPDATE: The lab teams continued preparing their methods, instruments, and workflow, facilitated by using ODP Leg 147 Hess Deep cores, samples, and data. Scientists and technical staff attended a presentation introducing some aspects of the Oman ophiolite. The Chief Scientists met with ship and drill crews to finalize the pre-drilling survey of sediment thickness as well as initial drilling/​coring operations. Our estimated time of arrival at Hess Deep Site HD-01B is 0100 h on 22 December.


Daily Science Report for 19 December 2012

LOCATION: In Transit to Hess Deep

SCIENCE UPDATE: The lab teams continued preparing their methods, instruments, and workflow, facilitated by using ODP Leg 147 Hess Deep cores, samples, and data. The USIO Operations Superintendent gave the final tour of the dynamic positioning system, rig floor, core tech shop, and mud pits. The Chief Scientists gave a presentation of the expedition science objectives and challenges to the ship’s crew. The science party received an orientation to wireline logging from the Logging Staff Scientist, followed by a visit to the logging tools. We continued our planning of the initial approach to seafloor beacon deployment and assessment of sediment thickness using a near bottom 3.5 kHz survey and the drill string. Our estimated time of arrival at Hess Deep Site HD-01B is 0600 h on 22 December.


Daily Science Report for 18 December 2012

LOCATION: In Transit to Hess Deep

SCIENCE UPDATE: Today, the lab teams continued preparing their methods, instruments, and workflow. We are using ODP Leg 147 Hess Deep cores, samples, and data to facilitate in this process. The USIO Operations Superintendent gave the science party an orientation to drilling and coring technology and Hess Deep operations and challenges. This was followed by a tour of the dynamic positioning system, rig floor, core tech shop, and mud pits. Later in the day, the Chief Scientists met with key staff to discuss initial seafloor beacon deployment and assessment of sediment thickness using near bottom 3.5 kHz survey and the drill string. A fire and boat safety drill was conducted in the morning followed by new scientists being given an orientation to the inside of a lifeboat by the Captain.


Daily Science Report for 17 December 2012

LOCATION: Commercial Wharf, Puntarenas, Costa Rica and In Transit to Hess Deep

SCIENCE UPDATE: Building on yesterday’s kick-off meetings, we continued with more in-depth scientist and technical staff lab team preparations. The igneous, metamorphic, and structure parts of the core description team split off to focus on their discipline-specific preparations. The paleomagnetism, geochemistry, and physical properties teams also met with their lab teams and identified existing Leg 147 sample residues to use for preparing their labs for new Hess Deep cores. We attended presentations on the local geology of the Hess Deep drill sites as well as the education and outreach plans for the expedition. We finished spooling the last of 9500 m of new core line on the aft drum, loaded the last two food shipments, and a backup logging tool. We departed the dock at 1818 h and began our ~4 day transit to Hess Deep Site HD-01B.


Daily Science Report for 16 December 2012

LOCATION: Commercial Wharf, Puntarenas, Costa Rica

SCIENCE UPDATE: We started the day with each scientist giving a short explanation of their primary research interests. Throughout the rest of the day we held kick-off meetings for each lab group, including an introduction to core description, physical properties, paleomagnetism, and geochemistry. The major activities for the crew included spooling new core line onto the aft drum, cleaning up after the bulk loading of drilling mud, loading of two containers of food, and storing/securing oncoming supplies. Anticipate departing Puntarenas late tomorrow afternoon.


Daily Science Report for 15 December 2012

LOCATION: Commercial Wharf, Puntarenas, Costa Rica

SCIENCE UPDATE: The Hess Deep scientists were given a presentation covering the expedition science expectations, procedures, deliverables, and obligations. Following this, the final tours of the laboratories and personal computer setup were conducted. Loading of the final drilling supplies and drilling mud were completed.


Daily Science Report for 14 December 2012

LOCATION: Commercial Wharf, Puntarenas, Costa Rica

SCIENCE UPDATE: Following Hess Deep scientist and technical staff introductions, the chief scientists presented an overview of the expedition science objectives. Scientists were given a tour of the laboratories and also underwent an orientation to shipboard computing and communications technology. Loading and storing of supplies continued – including drilling mud.


Daily Science Report for 13 December 2012

LOCATION: Commercial Wharf and at anchorage, Puntarenas, Costa Rica

SCIENCE UPDATE: Early in the morning, we moved to anchorage due to the arrival of two cruise ships. Key USIO staff met with the chief scientists to plan initial Expedition 345 operations. Expedition 345 scientists boarded the ship via water taxi and then underwent safety orientation. In the evening the cruise ships had departed, so we were able to return to the dock and resume normal port call activities that included loading of arriving shipments and preparing for tomorrows loading of drilling mud.


Daily Science Report for 12 December 2012

LOCATION: Commercial Wharf, Puntarenas, Costa Rica

SCIENCE UPDATE: Today, we continued with port call activities. This included ships crew change, offloading of temperature-controlled sample shipments and all departing sea freight. We loaded three containers of freight for the ships crew and a hydraulic pump required to repair a crane. Tomorrow, we will have to move the ship to anchorage for the entire day due to cruise ships. Scientists are scheduled to board tomorrow (13 December) via water taxi.


Daily Science Report for 11 December 2012

LOCATION: Commercial Wharf, Puntarenas, Costa Rica

SCIENCE UPDATE: The IODP Hess Deep Expedition (345) officially began when the ship arrived Puntarenas at 0500 h. We started our port call activities including customs and immigration, IODP-USIO crew change and crossover, boarding of chief scientists, and loading of airfreight. Due to shared dock space with cruise ships, all offloading/​loading activities are limited to evening hours. Scientists are scheduled to board the ship on 13 December.




About the IODP JRSO | Expeditions | Data & core samples | Tools & laboratories | Publications | Education |
Newsroom | Meetings & port calls | Employment | Contact us | Search | Site map | People | Intranet | Home

Accesibility Statement | Plug-in download | Privacy statement | ODP

For comments or questions: webmaster@iodp.tamu.edu

Copyright IODP JRSO

Modified on Saturday, 02-Feb-2013 18:56:58 CST.